# Hacking Task Confounder in Meta-Learning

### v1

@@ Introduction
where $\mathrm{z}_{i,j}^{s}$ is calculated the same as Eq.6. The objective of the outer loop mentioned in Eq.2 becomes:
@@
here, $\mathrm{z}_{i,j}^{q}$ is calculated as mentioned in Eq.7. Furthermore, in the second step, with $h$ and $g$ held fixed, we optimize $\Xi$ and $f_{gr}$ using Eq.6 and Eq.7.
In conclusion, MetaCRL, the meta-learning causal representation method we proposed, is a plug-and-play learner. By incorporating the causal invariant-based optimization mechanism and the additional regularization term, we can effectively eliminate task confounders that leads to model degradation and improve generalization capability.
Experiment
We evaluate MetaCRL on various scenarios, including sinusoid regression, image classification, drug activity prediction, and pose prediction. Considering that MetaCRL addresses the “Task Confounder” problem to enhance generalization, we compare it with two generalization baselines, MetaMix(Yao et al. 2021) and Dropout-Bins(Jiang et al. 2022). We also assess its performance on several backbones, including MAML(Finn, Abbeel, and Levine 2017), ANIL(Raghu et al. 2019), MetaSGD(Li et al. 2017), and T-NET(Lee and Choi 2018), to demonstrate its compatibility. Furthermore, the ablation study and visualization highlight the robustness of our method. We determine the hyperparameters based on the results shown in Appendix. More details about datasets, baselines, implementation, and additional experimental results can be found in Appendices C-F.
Sinusoid Regression
Firstly, we evaluate the performance of our MetaCRL on a sinusoid regression problem. Following (Jiang et al. 2022), the data for each task is generated in the form of $A\sin w\cdot x+b+\epsilon$, where $A\in\left[0.1,5.0\right]$, $w\in\left[0.5,2.0\right]$, and $b\in\left[0,2\pi\right]$. We add Gaussian observation noise with $\mu=0$ and $\epsilon=0.3$ to each data point sampled from the target task. In this experiment, we set $\lambda_{1}$ and $\lambda_{2}$ to 0.4 and 0.2. We use the Mean Squared Error (MSE) as the evaluation metric.
The results are shown in Table 1. Our method achieves improvements compared to the baselines, resulting in an average MSE reduction of 0.034 and 0.013, respectively. MetaCRL demonstrates even more substantial improvements across four different backbones, achieving an MSE reduction of over 0.1. Furthermore, we introduce IFSL (Yue et al. 2020) and MR-MAML (Yin et al. 2019) that constructed based on SCM, but their effects are less pronounced. As expected, our method exhibits significant enhancements, showcasing its high compatibility.
@@
Table 1: Performance (MSE) comparison on the sinusoid regression problem. The best results are highlighted in bold.
Image Classification
Next, we proceed to conduct tests in image classification, utilizing two benchmark datasets, miniImagenet and Omniglot. Notably, we introduce a specialized dataset called “TC”, which comprises 50 groups of tasks identified as being affected by task confounders. We measure the performance in the ”TC” dataset by comparing the experimental results with the performance of bachbones. More details about this dataset are provided in Appendix C. In this experiment, we set $\lambda_{1}$ and $\lambda_{2}$ to 0.5 and 0.35, respectively. The evaluation metric employed here is the average accuracy.
The results are shown in Table 2. Across all datasets, our method consistently surpasses the SOTA baseline. Notably, for the third data group, our approach outperforms the other baselines by a significant margin. This indicates that our MetaCRL can achieve similar or even better generalization improvements than baselines do without the need for task-specific or general-label space augmentation, while also demonstrating a unique advantage in handling task confounders. MetaCRL continues to exhibit remarkable performance and adeptly eliminates task confounders.
@@
Table 2: Performance (accuracy $\pm$ 95% confidence interval) of image classification on (20-way 1-shot) Omniglot and (5-way 1-shot) miniImagenet. See Appendix F for full results.
Drug Activity Prediction
Following (Yao et al. 2021), we evaluate MetaCRL on the drug activity prediction task (Martin et al. 2019). The dataset is designed to forecast the activity of compounds on specific target proteins, encompassing a total of 4276 tasks. In this experiment, $\lambda_{1}$ and $\lambda_{2}$ are both set to 0.3, and the evaluation metric is the squared Pearson correlation coefficient ($R^{2}$), reflecting the correlation between predictions and the actual values for each task. We record both the mean and median $R^{2}$ values, along with the count of $R^{2}$ values exceeding 0.3, which stands as a reliable indicator in pharmacology.
The results are shown in Table 3. Across the diverse sets of data, our approach attains performance levels akin to the SOTA baseline across all tasks. Notably, we achieve a noteworthy enhancement of 3 in the reliability index $R^{2}>0.3$. This achievement underscores the effectiveness of our approach across disparate domains and the pervasive influence of task confounders. See Appendix F for full results.
@@
ANIL + Dropout-Bins 0.394 0.321 53 0.338 0.271 48 0.312 0.284 46 0.370 0.297 50
ANIL + Ours 0.401 0.339 57 0.341 0.277 49 0.312 0.291 47 0.366 0.301 51
Table 3: Performance comparison on drug activity prediction. “Mean”, “Mde.”, and “$>0.3$” are the mean, the median value of $R^{2}$, and the number of analyzes for $R^{2}>0.3$. Values in the table is based on the results in (Jiang et al. 2022).
Pose Prediction
Lastly, we undertake the fourth benchmark, focusing on pose prediction. This task is constructed using the Pascal 3D dataset (Xiang, Mottaghi, and Savarese 2014). We randomly select 50 objects for meta-training and 15 additional objects for meta-testing. The values of $\lambda_{1}$ and $\lambda_{2}$ are set to 0.3 and 0.2. The evaluation metric employed here is MSE.
The results are shown in Table 4. Our MetaCRL achieves best performance. Notably, drawing insights from the findings presented in (Yao et al. 2021), we posit that augmenting the dataset could yield more effective results in this scenario, potentially outperforming the reliance solely on meta-regularization techniques. Thus, our approach incorporates regularization terms and still manages to achieve enhanced performance, thereby affirming its efficacy.
@@
ANIL 6.921 $\pm$ 0.415 6.602 $\pm$ 0.385
ANIL + MetaMix 6.394 $\pm$ 0.385 6.097 $\pm$ 0.311
ANIL + Dropout-Bins 6.289 $\pm$ 0.416 6.064 $\pm$ 0.397
ANIL + Ours 6.287 $\pm$ 0.401 6.055 $\pm$ 0.339
@@
Table 4: Performance (MSE $\pm$ 95% confidence interval) comparison on pose prediction.
(a) Sinusoid regression
(b) pose prediction
(c) miniImagenet
(d) Omniglot
Figure 3: Ablation study of MetaCRL on 4 benchmarks. The backbone in this experiment is MAML.
Figure 4: Accuracy (%) of models trained with different batch size on miniImagenet.
Figure 4: Accuracy (%) of models trained with different batch size on miniImagenet.
Figure 5: Visualization of the similarity matrix for causal factors in meta-training.
Ablation Study
@@
We conduct ablation study to explore the impact of different regularization terms, that is $\mathcal{L}_{\rm{DM}}(\Xi)$, $\mathcal{L}_{\rm{DM}}(f_{gr})$, and their combination $\mathcal{L}_{\rm{DM}}(f_{gr},\Xi)$. We select two classification and two regression scenarios from the aforementioned experiments for evaluation. The results, as shown in Figure 3, indicate that the first two regularization terms promote the model in all data sets, and the improvement is the largest when combined. Moreover, combining the aforementioned results, despite eliminating the regularization terms, our work still significantly outperforms the backbones, illustrating the effectiveness of the causal mechanism.
Visualization
@@
To better evaluate the effect of MetaCRL, we select MAML as the backbone and visualize the following metrics: (i) accuracy under different batch sizes; and (ii) the similarity between causal factors. The former evaluates MetaCRL’s efficacy in ensuring causality, while the latter assesses the decoupling of causal factors. Figures 5 and 5 show visualizations for these two aspects, respectively. Figure 5 shows that the model’s performance doesn’t decrease as batch size increases, which indicates that MetaCRL effectively eliminates task confounders. Figure 5 demonstrates that the disentangling module successfully decouples causal factors.
Conclusion
@@
In this paper, we propose a novel problem called “Task Confounder” and present a method called MetaCRL to address its unique challenges. We begin by analyzing a counterintuitive performance degradation phenomenon with SCM, revealing spurious correlations between causal factors of the training tasks and the generic label space, called “Task Confounder”. Then, we devise MetaCRL, which consists of two modules: (i) the disentangling module that acquires causal factors; (ii) the causal module that ensures causality of the factors. It is a plug-and-play causal representation learner that can be easily introduced into the meta-learning framework to eliminate task confounders. Extensive experiments demonstrate the effectiveness of our approach. Our work uncovers a novel and significant issue in meta-learning, providing valuable insights for future research.

### v4

@@ 4.3 Overall Objective
where $\mathrm{z}_{i,j}^{s}$ is calculated the same as Eq.7. Subsequently, the objective of the outer loop mentioned in Eq.2 becomes:
@@
where $\mathrm{z}_{i,j}^{q}$ is calculated as mentioned in Eq.8. Next, in the second step, with the meta-learning model $f_{\theta}$ held fixed, we optimize $\Xi$ and $f_{gr}$ as mentioned in Eq.7 and Eq.8.
By incorporating the causal invariant-based optimization mechanism and the additional regularization term, we can effectively eliminate task confounders that lead to model degradation and improve generalization capability.
5 Experiments
In this section, we first evaluate MetaCRL on various scenarios, including sinusoid regression, image classification, drug activity prediction, and pose prediction in Subsections 5.1-5.4, respectively. Next, we conduct ablation studies and visualization in Subsections 5.5 and 5.6. Considering that MetaCRL is a plug-and-play method, we assess its performance on several meta-learning models, e.g., MAML Finn et al. (2017), ANIL Raghu et al. (2019), MetaSGD Li et al. (2017), and T-NET Lee and Choi (2018), and multiple causal-based baselines, e.g., IFSL  Yue et al. (2020), Meta-Trans Bengio et al. (2019), Meta-Aug Rajendran et al. (2020), and MR-MAML  Yin et al. (2019), to demonstrate its compatibility. Considering that MetaCRL addresses the “Task Confounder” problem to enhance generalization, we also compare it with the plug-and-play generalization baselines that are most relevant to our method, i.e., MetaMix Yao et al. (2021) and Dropout-Bins Jiang et al. (2022). We delay all the details of datasets, baselines, implementation details, and additional experimental results in Appendices C-F, respectively.
@@ 5 Experiments
Table 1: Performance (MSE) comparison on the sinusoid regression problem. “+ours” means integrating MetaCRL into the existing methods, and the best results are highlighted in bold.
@@
IFSL 88.51 $\pm$ 0.49 36.21 $\pm$ 1.62 $\setminus$
Meta-Trans 87.39 $\pm$ 0.51 35.19 $\pm$ 1.58 $\setminus$
Meta-Aug 89.77 $\pm$ 0.62 34.76 $\pm$ 1.52 $\setminus$
MR-MAML 89.28 $\pm$ 0.59 35.01 $\pm$ 1.60 $\setminus$
@@
Table 2: Performance (accuracy $\pm$ 95% confidence interval) on (20-way 1-shot) Omniglot and (5-way 1-shot) miniImagenet. The “+” and “-” indicate the performance changes, and the $\setminus$” denotes that the result is not reported. See Appendix F for full results.
@@
ANIL + Dropout-Bins 0.394 0.321 53 0.338 0.271 48 0.312 0.284 46 0.368 0.297 50
ANIL + Ours 0.401 0.339 57 0.341 0.277 49 0.312 0.291 48 0.371 0.305 53
Table 3: Performance comparison on drug activity prediction. “Mean”, “Med.”, and “$>0.3$” are the mean, the median value of $R^{2}$, and the number of analyzes for $R^{2}>0.3$. The best results are highlighted in bold.
@@
Table 4: Performance (MSE $\pm$ 95% confidence interval) comparison on pose prediction. More results are provided in Appendix F.
5.1 Sinusoid Regression
Firstly, we evaluate the performance of our MetaCRL on sinusoid regression. Following  Jiang et al. (2022), we conduct 480 tasks and the data for each task is generated in the form of $A\sin w\cdot x+b+\epsilon$, where $A\in\left[0.1,5.0\right]$, $w\in\left[0.5,2.0\right]$, and $b\in\left[0,2\pi\right]$. We add Gaussian observation noise with $\mu=0$ and $\epsilon=0.3$ to each data point sampled from the target task. In this experiment, we set $\lambda_{1}$ and $\lambda_{2}$ to 0.4 and 0.2. We use the Mean Squared Error (MSE) as the evaluation metric.
The results are shown in Table 1. Compared to the plug-and-play baselines, MetaCRL achieves improvements with an average MSE reduction of 0.034 and 0.013, respectively. MetaCRL also demonstrates significant improvements across all the meta-learning base models, with an MSE reduction of over 0.1. Compared to the causal-based baselines, adding MetaCRL to any meta-learning model can always achieve better performance. As expected, MetaCRL exhibits significant enhancements, showcasing its high compatibility.
5.2 Image Classification
Next, we conduct experiments on image classification, utilizing two benchmark datasets, i.e., miniImagenet and Omniglot. These two datasets contain 600 and 1623 tasks, respectively. We also introduce a specialized dataset called “TC”, which comprises 50 groups of tasks (300 tasks in total) identified as being affected by task confounders, i.e., tasks with negative knowledge transfer as mentioned in Subsection 3.2. More details are provided in Appendix C. In this experiment, we set $\lambda_{1}$ and $\lambda_{2}$ to 0.5 and 0.35, respectively. The evaluation metric employed here is the average accuracy.
The results are shown in Table 2. MetaCRL consistently surpasses the SOTA baselines across all datasets, indicating that it can achieve better generalization improvements than the baselines do without the need for task-specific or general-label space augmentation that the baselines need. Notably, on the “TC” dataset, MetaCRL outperforms the baselines by a significant margin, which demonstrates a unique advantage of MetaCRL in handling task confounders. In summary, MetaCRL continues to exhibit remarkable performance and adeptly eliminates task confounders.
5.3 Drug Activity Prediction
We also evaluate MetaCRL on drug activity prediction. pQSAR Martin et al. (2019) is a dataset designed to forecast the activity of compounds on specific target proteins, encompassing a total of 4276 tasks. We adopt the same settings as Yao et al. (2021) and divide the tasks into four groups. In this experiment, $\lambda_{1}$ and $\lambda_{2}$ are both set to 0.3, and the evaluation metric is the squared Pearson correlation coefficient ($R^{2}$), reflecting the correlation between predictions and the actual values for each task. We record both the mean and median $R^{2}$ values, along with the count of $R^{2}$ values exceeding 0.3, which stands as a reliable indicator in pharmacology.
The results are shown in Table 3. MetaCRL attains performance levels akin to the SOTA baselines across all four groups of data. Notably, we achieve a noteworthy enhancement of 3 in the reliability index $R^{2}>0.3$. The achievement of this scenario underscores the effectiveness of our MetaCRL across disparate domains and the pervasive influence of task confounders. See Appendix F for full results.
Figure 3: Ablation study, including (a) sinusoid regression, (b) pose prediction, (c) 5-way 1-shot miniImagenet, and (d) 20-way 1-shot Omniglot. The backbone is MAML. The red, blue, green, and orange bars represent the results of MetaCRL-$\mathcal{L}_{\rm{DM}}(f_{gr},\Xi)$, MetaCRL-$\mathcal{L}_{\rm{DM}}(\Xi)$, MetaCRL-$\mathcal{L}_{\rm{DM}}(f_{gr})$, and MetaCRL.
5.4 Pose Prediction
Lastly, we undertake the fourth benchmark, focusing on pose prediction. This evaluation is constructed using the Pascal 3D dataset  Xiang et al. (2014). We randomly select 50 objects for meta-training and 15 additional objects for meta-testing. In this experiment, the values of $\lambda_{1}$ and $\lambda_{2}$ are set to 0.3 and 0.2, while the evaluation metric employed here is MSE.
The results are shown in Table 4. MetaCRL achieves the best performance. Notably, drawing insights from the findings presented in  Yao et al. (2021), we posit that augmenting the dataset could yield more effective results in this scenario, potentially outperforming the reliance solely on meta-regularization techniques. MetaCRL incorporates regularization terms instead of data augmentation and still manages to achieve enhanced performance, thereby affirming its efficacy.
5.5 Ablation Study
@@
We conduct ablation studies to explore the impact of different regularization terms, that is $\mathcal{L}_{\rm{DM}}(\Xi)$, $\mathcal{L}_{\rm{DM}}(f_{gr})$, and their combination $\mathcal{L}_{\rm{DM}}(f_{gr},\Xi)$ in Eq.6. We select both classification and regression scenarios, including four benchmark datasets. Figure 3 shows the results that $\mathcal{L}_{\rm{DM}}(\Xi)$ and $\mathcal{L}_{\rm{DM}}(f_{gr})$ promote the model in all datasets, and the improvement is the largest when combined. Moreover, despite eliminating the regularization terms, MetaCRL still significantly outperforms the base models, illustrating the effectiveness of the causal module. We also construct ablation studies targeting the accuracy of extracting task-specific causal factors and model efficiency (See Appendix F for details).
Figure 4: Knowledge transference after using MetaCRL.
Figure 5: Visualization of the similarity between causal factors.
5.6 Visualization
@@
To better evaluate the effect of MetaCRL, we visualize (i) knowledge transfer after using MetaCRL; and (ii) the similarity between causal factors. The former evaluates MetaCRL’s efficacy in ensuring causality and avoiding negative knowledge transfer caused by task confounders, which use the same settings as in Subsection 3.2. The latter assesses the decoupling of causal factors using cosine similarity. Figures 5 and 5 show visualizations for these two aspects, respectively. Figure 5 shows that there are almost no training tasks that lead to negative knowledge transfer with fewer iterations than Figure 1, which indicates that MetaCRL effectively eliminates task confounders. Figure 5 shows that the similarity scores between different causal factors are very low, illustrating that the disentangling module successfully decouples causal factors. More details are provided in Appendix F.
6 Conclusion
@@
In this paper, we discover a valuable problem called “Task Confounder”, and propose a novel method called MetaCRL to address its unique challenges. We begin by analyzing a counterintuitive negative knowledge transfer phenomenon with SCM, revealing spurious correlations between causal factors of the training tasks and the label space, i.e., “Task Confounder”. Then, we propose MetaCRL, which consists of two modules: (i) a disentangling module that acquires generating factors and eliminates task confounders; and (ii) a causal module that ensures causality of the obtained generating factors. It is a plug-and-play causal representation learner that can be applied to any meta-learning baseline. Extensive experiments demonstrate the effectiveness and robustness of MetaCRL. Our work uncovers a novel and significant issue in ML, providing valuable insights for future research.
Acknowledgements
The authors would like to thank the anonymous reviewers for their valuable comments. This work is supported in part by the Postdoctoral Fellowship Program of CPSF No. GZB20230790, the China Postdoctoral Science Foundation No. 2023M743639, and the Special Research Assistant Fund, Chinese Academy of Sciences No. E3YD590101. The Appendix is provided in https://arxiv.org/abs/2312.05771.
Contribution Statement
Jingyao Wang and Yi Ren made equal contributions. All the authors participated in designing research, performing research, analyzing data, and writing the paper.